OpenStudy (anonymous):

y=x-4/x^2-16 what are the asymptotes ?

OpenStudy (anonymous):

factor and cancel first

OpenStudy (anonymous):

if you factor its y=1/x-4

OpenStudy (anonymous):

no

OpenStudy (anonymous):

when you cancel the \(x-4\) you get \[\frac{1}{x+4}\]

OpenStudy (anonymous):

which of the following is not an asymptote x = 4 x = -4 y = 0 Both x = 4 and y = 0 All of these are asymptotes.

OpenStudy (anonymous):

??????

OpenStudy (anonymous):

one you get \(\frac{1}{x+4}\) find the asymptote by setting \(x+4=0\) and solve for \(x\) (in your head)

OpenStudy (anonymous):

x=-4

OpenStudy (anonymous):

you get \(x=-4\) IS an asymptote, the others are not

OpenStudy (anonymous):

so x=4 and the rest wouldn't be asymptotes only x=-4?